EXPECTED AND REQUIRED RATES OF RETURN Assume that the risk-free rate is 4% and the market risk premium is 6%. What is the required return for the overall stock market? Round your answer to two decimal places. % What is the required rate of return on a stock with a beta of 1.2? Round your answer to two decimal places. %
Answer #1

market risk premium=market rate-risk free rate

1.market rate=(4+6)=10%

2.required return= risk-free rate +Beta*(MArket rate- risk-free rate )

=4+1.2*(10-4)

=11.2%